# Simulation of charged nanotubes self-assembly during evaporation of a sessile droplet on a substrate

### v1

@@ Appendix B Coefficient of friction
Let us make a rough estimate of the viscous friction coefficient $k_{f}$ for the nanotubes under consideration. To do this, we need to estimate the Reynolds number, $\mathrm{Re}=dv_{c}\rho/\eta\approx 10^{-6}$, where the characteristic flow velocity is $v_{c}\approx$ 10 µm/s [26]. As the first approximation, we use a linear dependence on nanotube velocity $F_{d}=k_{f}v_{c}$. The drag force is expressed as $F_{d}=F_{d\mathrm{Lamb}}\left(1-0.87\lambda^{2}+0.514\left(1-\exp(-\mathrm{Re})\right)\lambda^{3}\right)$, where $\lambda=\left(0.5-\gamma-\ln(\mathrm{Re}/8)\right)^{-1}$, $F_{d\mathrm{Lamb}}=4\pi\eta v_{c}\lambda l$, and the Euler constant is equal to $\gamma\approx 0.577$ [27, 28]. As a result, we obtain $k_{f}=F_{d}/v_{c}\approx 5\times 10^{-10}$ kg/s. When some nanotube rotates in a medium, the medium exerts resistance. Assuming that the drag force is applied to one end of the nanotube, the magnitude of the viscous torque is approximated by the formula $M_{\mathrm{vis}}=-k_{f}\,\omega\,(l/2)^{2}$.

### v2

@@ Appendix B Coefficient of friction
Let us make a rough estimate of the viscous friction coefficient $k_{f}$ for the nanotubes under consideration. To do this, we need to estimate the Reynolds number, $\mathrm{Re}=dv_{c}\rho/\eta_{0}\approx 10^{-6}$, where the characteristic flow velocity is $v_{c}\approx$ 10 µm/s [16]. As the first approximation, we use a linear dependence on nanotube velocity $F_{d}=k_{f}v_{c}$. The drag force is expressed as $F_{d}=F_{d\mathrm{Lamb}}\left(1-0.87\lambda^{2}+0.514\left(1-\exp(-\mathrm{Re})\right)\lambda^{3}\right)$, where $\lambda=\left(0.5-\gamma-\ln(\mathrm{Re}/8)\right)^{-1}$, $F_{d\mathrm{Lamb}}=4\pi\eta_{0}v_{c}\lambda l$, and the Euler constant is equal to $\gamma\approx 0.577$ [29, 30]. As a result, we obtain $k_{f}=F_{d}/v_{c}\approx 5\times 10^{-10}$ kg/s. When some nanotube rotates in a medium, the medium exerts resistance. Assuming that the drag force is applied to one end of the nanotube, the magnitude of the viscous torque is approximated by the formula $M_{\mathrm{vis}}=-k_{f}\,\omega\,(l/2)^{2}$.